OpenStudy (anonymous):

Find the distance d(p1, p2) between the points p1, p2. p1=(4,-4); p2=(5,1)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

I got square root of 26 so far.

OpenStudy (anonymous):

if I simplify that it goes into 2 * 13?

OpenStudy (anonymous):

is 2 square root of 13 the answer?

hartnn (hartnn):

sqrt 26 is the correct answer u can't simplify that to 2 square root of 13

OpenStudy (mimi_x3):

I think the formula should be \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] or is it the same thing?

hartnn (hartnn):

same thing

OpenStudy (anonymous):

how come this answer I got right it was 2 square root of 17 the problem was p1(4,-3) pr(2,5)

OpenStudy (anonymous):

it came out to be square root of 68 that couldn't be simplified right?

hartnn (hartnn):

yeah, \(\huge \sqrt {68}= \sqrt {4*17}= \sqrt 4 \sqrt 17 =2\sqrt17\)

OpenStudy (anonymous):

oh ok I see

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcomes ^_^